# The electronic structure and disorder effect of La3Ni2O7 superconductor

### v1

@@ Abstract
Determining the electronic structure of La3Ni2O7 is an essential step towards uncovering their superconducting mechanism. It is widely believed that the bilayer apical oxygens play an important role in the bilayer La3Ni2O7 electronic structure. Applying the hybrid exchange-correlation functionals, we obtain a more accurate electronic structure of La3Ni2O7 at its high-pressure phase, where the binding $d_{z^{2}}$ band is below the Fermi level owing to apical oxygen. The symmetry properties of this electronic structure and its corresponding tight-binding model are further analyzed. We find the antisymmetric part is highly entangled leading to a minimal nearly degenerate two-orbital model. Then, the apical oxygen vacancies effect is studied using the dynamical cluster approximation. This disorder effect strongly destroys the antisymmetric $\beta$ Fermi surface leading to the possible disappearance of superconductivity.
@@
In this work, we focus on the La3Ni2O7 electronic structure at the superconducting phase. We start from the electronic structure of La3Ni2O7 at its $Fmmm$ structure. We apply the density functional theory (DFT) calculation with hybrid functional, which more accurately treats the Coulomb potential. This method has given a faithful description for the ambient pressure La3Ni2O7 electronic and magnetic structures Wang et al. (2024c); Chen et al. (2024b); Yang et al. (2024). The result is consistent with our previous findings, where the $d_{z^{2}}$ bonding band is always below the Fermi level. Then, the influence of apical oxygen and the random disorder effect of apical oxygen is explored by the dynamical cluster approximation (DCA) Hettler et al. (2000); Jarrell and Krishnamurthy (2001); Maier et al. (2005).
@@
Figure 3: The calculated spectrum function $A({\textit{{k}}},\omega)$ using the lattice Green’s function with the DCA self-energy $\Sigma({\textit{{K}}},\omega)$ along the high-symmetry lines and high-symmetry points for various concentrations of the apical oxygen vacancies $\delta$. Here $\omega$ is measured from the chemical potential.
@@
which reflects the disorder effect due to the random oxygen vacancies.
Figure 4: The calculated spectrum function at the chemical potential $A({\textit{{k}}},0)$ using the lattice Green’s function with the DCA self-energy $\Sigma({\textit{{K}}},0)$ in the whole Brillouin zone for various concentrations of the apical oxygen vacancies $\delta$. As the vacancies increase, the $\beta$-FS becomes more vague while the $\alpha$-FS remains sharp.
@@
In this section, we present the fit of the TB model to the HSE band structure and GGA-PBE band structure of $a=3.71\AA$ obtained from the DFT calculations, as shown in Fig. S2. It can be seen that TB and DFT are in excellent agreement, indicating that our TB model provides an accurate representation of the $e_{g}$ orbital bands. The TB model of $a=3.71\AA$ thin film is also listed in Table. S1.
@@
$t_{1}^{x}$ $t_{1}^{z}$ $t_{2}^{x}$ $t_{2}^{z}$ $t_{3}^{xz}$
-0.4907 -0.0757 0.0868 -0.0123 0.2007
$t_{\bot}^{x}$ $t_{\bot}^{z}$ $t_{4}^{xz}$ $\epsilon^{x}$ $\epsilon^{z}$
0.0107 -0.6134 -0.0267 0.9412 0.1497
sector $t_{1}^{x}$ $t_{1}^{z}$ $t_{2}^{x}$ $t_{2}^{z}$ $t_{3}^{xz}$ $\epsilon^{x}$ $\epsilon^{z}$
$H_{S}$0.4907 -0.0757 0.0868 -0.0123 0.174 0.9519 -0.4637
$H_{A}$0.4907 -0.0757 0.0868 -0.0123 0.2274 0.9305 0.7631

### v2

@@ Abstract
Determining the electronic structure of La3Ni2O7 is an essential step towards uncovering their superconducting mechanism. It is widely believed that the bilayer apical oxygens play an important role in the bilayer La3Ni2O7 electronic structure. Applying the hybrid exchange-correlation functionals, we obtain a more accurate electronic structure of La3Ni2O7 at its high-pressure phase, where the bonding $d_{z^{2}}$ band is below the Fermi level owing to apical oxygen. The symmetry properties of this electronic structure and its corresponding tight-binding model are further analyzed. We find the antisymmetric part is highly entangled leading to a minimal nearly degenerate two-orbital model. Then, the apical oxygen vacancies effect is studied using the dynamical cluster approximation. This disorder effect strongly destroys the antisymmetric $\beta$ Fermi surface leading to the possible disappearance of superconductivity.
Electronic structure; Oxygen vacancies disorder; Dynamical cluster approximation; Bilayer superconducting nickelate
pacs:
71.23.-k
@@
In this work, we focus on the La3Ni2O7 electronic structure at the superconducting phase. The latest experimental results indicate that when the pressure is approximately 18 GPa, the crystal structure adopts the $Fmmm$ space group, at which point the superconducting transition temperature reaches its maximum. When the pressure exceeds 40 GPa, the space group of the crystal structure transitions to $I4/mmm$ li2024pressure. Therefore, we can start from the electronic structure of La3Ni2O7 at its $Fmmm$ structure. We apply the density functional theory (DFT) calculation with hybrid functional, which more accurately treats the Coulomb potential. This method has given a faithful description for the ambient pressure La3Ni2O7 electronic and magnetic structures Wang et al. (2024c); Chen et al. (2024b); Yang et al. (2024). The result is consistent with our previous findings, where the $d_{z^{2}}$ bonding band is always below the Fermi level. Then, the influence of apical oxygen and the random disorder effect of apical oxygen is explored by the dynamical cluster approximation (DCA) Hettler et al. (2000); Jarrell and Krishnamurthy (2001); Maier et al. (2005).
@@
Figure 3: The calculated spectrum function $A({\textit{{k}}},\omega)$ using the lattice Green’s function with the DCA self-energy $\Sigma({\textit{{K}}},\omega)$ along the high-symmetry lines and high-symmetry points for various concentrations of the apical oxygen vacancies $\delta$. Here $\omega$ is measured from the chemical potential. The impurity potential is modeled by assuming the apical oxygen vacancy completely removes the interlayer intracell coupling $t_{\perp}^{x}$ and $t_{\perp}^{z}$ and the calculation is performed at zero temperature.
@@
which reflects the disorder effect due to the random oxygen vacancies. The DCA calculation is performed at zero temperature.
Figure 4: The calculated spectrum function at the chemical potential $A({\textit{{k}}},0)$ using the lattice Green’s function with the DCA self-energy $\Sigma({\textit{{K}}},0)$ in the whole Brillouin zone for various concentrations of the apical oxygen vacancies $\delta$. As the vacancies increase, the $\beta$-FS becomes more vague while the $\alpha$-FS remains sharp. The impurity potential is modeled by assuming the apical oxygen vacancy completely removes the interlayer intracell coupling $t_{\perp}^{x}$ and $t_{\perp}^{z}$ and the calculation is performed at zero temperature.
@@
Although direct ARPES measurements on high-pressure phases are currently unavailable, two indirect approaches can still be employed to verify the validity of our proposed scenario. First, for La3Ni2O7 thin films grown on SrLaAlO4 (SLAO) substrates—which exhibit superconductivity under ambient pressure Ko et al. (2024); Zhou et al. (2024)—we can directly measure the evolution of the $\beta$-Fermi surface using ARPES when oxygen vacancies are introduced. Second, the superconducting pairing symmetry can be experimentally determined. Prior studies have debated this extensively, proposing both $s^{\pm}$-wave and $d$-wave possibilities Luo et al. (2023); Yang et al. (2023a); Gu et al. (2023); Liu et al. (2023); Lu et al. (2024); Zhang et al. (2024b); Ryee et al. (2023); Jiang et al. (2024). If experiments confirm a $d$-wave symmetry, this would strongly suggest that the $\beta$-band dominates the superconducting pairing, thereby supporting the conclusions of our work.
@@
In this section, we present the fit of the TB model to the HSE band structure and GGA-PBE band structure of $a=3.71\AA$ obtained from the DFT calculations, as shown in Fig. S2. It can be seen that TB and DFT are in excellent agreement, indicating that our TB model provides an accurate representation of the $e_{g}$ orbital bands. The TB model of $a=3.71\AA$ thin film is also listed in Table. S1. The use of the same number of parameters as in the original model in the main text cannot achieve a satisfactory fit between the TB and DFT band structures. To achieve a better fit, we incorporated additional longer-range hopping parameters into the TB model, including: intralayer same-orbital third-nearest-neighbor hopping parameter $t_{5}$, intralayer inter-orbital third-nearest-neighbor hopping parameter $t_{6}$, interlayer inter-orbital third-nearest-neighbor hopping parameter $t_{7}$ (note that the inter-orbital second-nearest-neighbor hopping parameters are zero due to symmetry constraints), and interlayer same-orbital nearest-neighbor hopping parameter $t_{\bot 1}$ (with the interlayer on-site hopping parameter denoted as $t_{\bot}$). Specifically, the $T_{\textbf{k}}$, $V_{\textbf{k}}$ and $V_{\textbf{k}}$ are written as:
$\begin{split}T_{\textbf{k}}^{x/z}&=t_{1}^{x/z}\gamma_{k}+t_{2}^{x/z}\alpha_{k}+t_{5}^{x/z}\delta_{k},\\ V_{\textbf{k}}&=t_{3}^{xz}\beta_{k}+t_{6}^{xz}\theta_{k},\\ V_{\textbf{k}}^{\prime}&=t_{4}^{xz}\beta_{k}+t_{7}^{xz}\theta_{k},\end{split}$ (22)
where $\delta_{k}=2(\cos 2k_{x}+\cos 2k_{y})$, $\theta_{k}=2(\cos 2k_{x}-\cos 2k_{y})$. The matrix element originally denoted as $t_{\bot}^{x/z}$ should also be modified to $t_{\bot}^{x/z}+t_{\bot 1}^{x/z}\gamma_{k}$.
@@
$t_{1}^{x}$ $t_{1}^{z}$ $t_{2}^{x}$ $t_{2}^{z}$ $t_{3}^{xz}$ $t_{4}^{xz}$ $t_{5}^{x}$ $t_{5}^{z}$
-0.4907 -0.0757 0.0868 -0.0123 0.2007 -0.0267 -0.0588 -0.0148
$t_{6}^{xz}$ $t_{7}^{xz}$ $t_{\bot}^{x}$ $t_{\bot}^{z}$ $t_{\bot 1}^{x}$ $t_{\bot 1}^{z}$ $\epsilon^{x}$ $\epsilon^{z}$
0.0274 0.0007 0.0107 -0.6134 0.0039 0.0021 1.0112 0.2197
sector $t_{1}^{x}$ $t_{1}^{z}$ $t_{2}^{x}$ $t_{2}^{z}$ $t_{3}^{xz}$ $t_{5}^{x}$ $t_{5}^{z}$ $t_{6}^{xz}$ $\epsilon^{x}$ $\epsilon^{z}$
$H_{S}$0.4868 -0.0736 0.0868 -0.0123 0.174 -0.0588 -0.0148 0.0281 1.0219 -0.3937
$H_{A}$0.4946 -0.0778 0.0868 -0.0123 0.2274 -0.0588 -0.0148 0.0267 1.0012 0.8331